A retailer purchased 400 hats at $15 each. He sold 250 at the regular price of $40, 100 at
20% off the regular price and the remaining hats were cleared out at $15 each. The
retailer's overhead is 35% of the cost. What is the total operating profit (or loss) on the sale
of all the hats?
1
Expert's answer
2020-06-01T18:10:02-0400
"C=\$15\cdot400(1+{35\over 100})=\$8100"

"R=\$40\cdot250+\$40({100-20\over 100})\cdot 100+\$15\cdot(400-250-100)=""=\$13950"

"P=R-C=\$13950-\$8100=\$5850"

The total operating profit on the sale of all the hats is "\$5850."